You are making a drink that requires 12 ounces of a 33-proof whiskey. Unfortunately, all you have on hand are 24-proof and 36-proof whiskey. How much of each type of whiskey do you need to mix to create 12 ounces of a 33-proof whiskey? The proof of an alcoholic beverage indicates the percentage of alcohol (ethanol) it contains. A 24-proof whiskey, for example, contains 48 percent alcohol by volume. You need to mix ? oz. of 24-proof with ? oz. of 36-proof to produce 12 ounces of 33-proof whiskey.
Answer by Alan3354(69443)

You are making a drink that requires 12 ounces of a 33-proof whiskey. Unfortunately, all you have on hand are 24-proof and 36-proof whiskey. How much of each type of whiskey do you need to mix to create 12 ounces of a 33-proof whiskey?
s = 36 proof amount
w = 24 proof amount
----
s + w = 12 (total liquid)
36s + 24w = 12*33 (total proof)
-----------------------
The proof of an alcoholic beverage indicates the percentage of alcohol (ethanol) it contains. A 24-proof whiskey, for example, contains 48 percent alcohol by volume.
It's the other way around. 24 proof is 12%.
---
You need to mix ? oz. of 24-proof with ? oz. of 36-proof to produce 12 ounces of 33-proof whiskey.
-----------
Solve for s & w.